# Soban Nasir Lone

### v1

@@ Abstract
We introduce a novel deep operator network (DeepONet) framework that incorporates generalised variational inference (GVI) using Rényi’s $\alpha$-divergence to learn complex operators while quantifying uncertainty. By incorporating Bayesian neural networks as the building blocks for the branch and trunk networks, our framework endows DeepONet with uncertainty quantification. The use of Rényi’s $\alpha$-divergence, instead of the Kullback-Leibler divergence (KLD), commonly used in standard variational inference, mitigates issues related to prior misspecification that are prevalent in Variational Bayesian DeepONets. This approach offers enhanced flexibility and robustness. We demonstrate that modifying the variational objective function yields superior results in terms of minimising the mean squared error and improving the negative log-likelihood on the test set. Our framework’s efficacy is validated across various mechanical systems, where it outperforms both deterministic and standard KLD-based VI DeepONets in predictive accuracy and uncertainty quantification. The hyperparameter $\alpha$, which controls the degree of robustness, can be tuned to optimise performance for specific problems. We apply this approach to a range of mechanics problems, including gravity pendulum, advection-diffusion, and diffusion-reaction systems. Our findings underscore the potential of $\alpha$-VI DeepONet to advance the field of data-driven operator learning and its applications in engineering and scientific domains.
Keywords Prior-Robust Bayesian Inference  $\cdot$ Alpha Divergence  $\cdot$ Variational Bayes DeepONet  $\cdot$ Bayesian Neural Networks  $\cdot$ Uncertainty quantification  $\cdot$ PDE Surrogate
@@ 1 Introduction
As scientific machine learning methodologies take centre stage across diverse disciplines in science and engineering, there is an increased interest in adopting data-driven methods to analyse, emulate, and optimise complex physical systems. The behaviour of such systems is often described by laws expressed as systems of ordinary differential equations (ODEs) and partial differential equations (PDEs) [1]. A classical task then involves the use of analytical or computational tools to solve such equations across a range of scenarios, e.g., different domain geometries, input parameters, and initial and boundary conditions (IBCs). Solving these so-called parametric PDE problems requires learning the solution operator that maps variable input entities to the corresponding solution of the underlying PDE system. Tackling this task using traditional tools (e.g., finite element methods [2]) bears a formidable cost, as independent simulations need to be performed for every different domain geometry, input parameters, or IBCs. Driven by this challenge, a growing field of neural operators for solving parametric PDEs has come forth recently.
@@
Learning to solve PDEs is closely related to operator learning. Instead of learning to solve a specific PDE, it can be advantageous to learn the operator that maps a functional parameter of the PDE (such as initial values, boundary conditions, force fields, or material parameters) to the solution associated with the given parameter. Neural operators exemplify this approach by learning to solve entire classes of PDEs simultaneously. These specially designed deep neural networks can represent the solution map of parametric PDEs in a discretisation-invariant manner, meaning the model can be queried at any arbitrary output location. Unlike neural networks that map between two finite-dimensional vector spaces, neural operators can represent mappings between spaces with infinite dimensions [3]. Popular operator learning frameworks include deep operator networks (DeepONets) [4] and Fourier neural operator (FNO) [5]. This article focuses on DeepONet, which has been applied to a wide range of problems, including solid and fluid mechanics [6], reliability analysis [7], heat transfer [8], and fracture mechanics [9]. De et al. [10] utilised DeepONets for modelling uncertain and partially unknown systems. Cai et al. [11] applied DeepONets to model field variables across multiple scales in multi-physics problems while He et al. [12] employed DeepONets to predict full-field, nonlinear elastic–plastic stress responses in complex geometries. Other notable applications of DeepONets include weather forecasting [13], reduced-order modelling [14], finance [15], and others [16, 17].
@@
While DeepONets excel at solving differential equations (DEs), a crucial aspect often overlooked is the uncertainty quantification of the predictions. Traditional DeepONet architectures have been found to produce over-confident predictions, implying they are poorly calibrated [18]. This means they might underestimate the true range of possible outcomes, potentially leading to unreliable results. Additionally, they do not quantify the inherent uncertainties associated with their predictions. In engineering applications like aerodynamics or structural analysis, uncertainty quantification is paramount. By quantifying uncertainties in predictions related to phenomena such as stress distribution, engineers can make informed decisions about the risk and reliability of their designs.
@@
Several attempts have been made to incorporate uncertainty quantification within the DeepONet architecture. However, existing approaches face some limitations. Lin et al. [19] proposed a Bayesian DeepONet based on replica-exchange stochastic gradient Langevin diffusion [20] that considers the standard deviation of the output to be known, which may not always be realistic. Additionally, training the replicas in this approach is computationally expensive. Yang et al. [21] used randomised priors and trained an ensemble of deterministic models, which can also be computationally demanding depending upon the size of the ensemble. While the method of ensembles offers some degree of uncertainty estimation through multiple model predictions, it does not adhere to the more principled framework that is provided by Bayes’ rule and hence cannot be considered truly Bayesian.
@@
Variational inference (VI)-based Bayesian methods offer a promising framework for incorporating uncertainty quantification into deep learning models. This approach has given rise to Bayesian neural networks (BNNs) [22, 23]. In BNNs, the parameters of the neural network are treated as random variables, which allow the uncertainty in the parameters to propagate through the network, ultimately reflecting the uncertainty in the model predictions. The variational inference approach is particularly advantageous for neural networks with a large number of parameters, as traditional Markov chain Monte Carlo (MCMC)-based methods become computationally expensive to use. Variational inference, on the other hand, provides a scalable alternative for approximating the posterior distribution of the parameters in BNN.
@@
It seems natural to extend the idea of uncertainty quantification in BNNs to that of a Bayesian DeepONet. Building on this idea, Garg et al. [24] proposed VB-DeepONet that uses BNNs as building blocks for the DeepONet architecture. However, VB-DeepONet inherits certain challenges associated with BNNs. Their approach employs a fully factorised standard Gaussian distribution as the prior over the DeepONet model parameters. Such a prior might not accurately reflect the true distribution of these parameters, leading to a sub-optimal approximation of the posterior distribution. As indicated in Li et al. [25], Knoblauch et al. [26], and Wenzel et al. [27], such a misspecified prior can hinder the overall effectiveness of variational inference, where the approximated posterior concentrates around a single point leading to overconfident predictions and underestimated uncertainty [28]. This hinders the model’s ability to effectively quantify uncertainties.
@@
We propose to address these limitations in VB-DeepONet by introducing prior-robust variational inference for DeepONets. Standard VI relies on the Kullback-Leibler divergence (KLD), which is sensitive to prior selection, especially when the number of parameters is large. Alternative divergence measures exist that offer robustness to misspecified priors. Knoblauch et al. [26] introduced the Generalised Variational Inference (GVI) framework that allows for defining divergence metrics beyond KLD in the context of BNNs. GVI incorporates a hyperparameter that controls the degree of robustness to prior misspecification. This leads to posteriors that are less influenced by priors which deviate significantly from the observed data [26].
@@
In this paper, we extend the GVI framework to DeepONets to achieve greater robustness to prior misspecification. In particular, we propose the use of Rényi’s $\alpha$-divergence [29] as a robust alternative to KLD within the VI framework. This novel operator learning approach using GVI leads to an improved approximated posterior distribution, demonstrably enhancing both prediction accuracy and the quality of uncertainty estimates as evidenced by log-likelihood values on unseen test data.
The rest of the article will provide background on deterministic DeepONets, followed by details of the proposed variational Bayesian framework and four numerical examples highlighting the efficacy of our approach. We conclude by discussing the results of different hyperparameter settings and recovering standard KLD-VI results as a special case.
@@ 2 Background on DeepONets
Operator networks were first introduced by Chen and Chen [30], where they considered shallow networks with a single hidden layer. Subsequently, this concept was significantly developed with deep neural network architectures, leading to the creation of DeepONets [4]. To define operators, it is useful to consider two different classes of functions residing in two separate Banach spaces, $\mathcal{A}$ and $\mathcal{S}$, respectively. One class of functions, $a(\bm{{y}})\in\mathcal{A}$, with domain in $\Omega_{Y}\in\mathbb{R}^{D}$, while the other class of functions, $s(\bm{{y}})\in\mathcal{S}$, also having domain in $\Omega_{Y}$. An operator $\mathcal{G}$ can be defined as a mapping between these two function spaces, $\mathcal{G}:\mathcal{A}\rightarrow\mathcal{S}$ such that $s(\bm{{y}})=\mathcal{G}(a(\bm{{y}}))$.
@@ 2.1 Deterministic DeepONets
Here, $\mathcal{F}$ is a nonlinear differential operator, $x$ is the location in space, $t$ is the time; $s$ denotes the solution of the differential equation, and $u$ denotes the external source function. Furthermore, $\mathcal{B}(s)(x,t)=s_{b}(x)$ and $\mathcal{I}(s)(x,0)=s_{0}(x)$ define the boundary and initial conditions, respectively. The true solution operator is denoted by $\mathcal{G}(u)(x,t)$. To keep the notations compact, we denote the tuple $(x,t)$ by $\bm{{y}}$.
DeepONets [4] are designed to learn an approximate operator ${\mathcal{G}}_{\theta}$ parameterised by a deep neural network with parameter vector $\bm{{\theta}}$.
@@
The architecture for the DeepONet consists of two main sub-networks: a branch neural network and a trunk neural network (see Fig. 1).
@@
$\begin{split}\hat{\bm{{\theta}}}&=\operatorname*{argmin}_{\bm{{\theta}}}\frac{1}{N_{1}N_{2}}\sum_{i=1}^{N_{1}}\sum_{k=1}^{N_{2}}\left(\mathcal{G}_{\bm{{\theta}}}\left(\bm{{a}}^{(i)}\right)(\bm{{y}}_{k})-s^{(i)}\left(\bm{{y}}_{k}\right)\right)^{2}\\ &=\frac{1}{N_{1}N_{2}}\sum_{i=1}^{N_{1}}\sum_{k=1}^{N_{2}}\left(b_{0}+\sum_{p=1}^{P}b_{p}\left(\bm{{a}}^{(i)};\bm{{\theta}}_{B}\right)\;\psi_{p}\left(\bm{{y}}_{k};\bm{{\theta}}_{T}\right)-s^{(i)}\left(\bm{{y}}_{k}\right)\right)^{2}.\end{split}$ (4)
The optimisation problem can be solved using stochastic gradient descent [31] to obtain the optimal parameter values for the DeepONet. This approach is referred to as D-DeepONet in the rest of the article, where “D” stands for deterministic.
@@ 2.2 Variational Bayesian DeepONets (VB-DeepONet)
VB-DeepONet [24] addresses the limitations of D-DeepONet by incorporating layers of Bayesian neural networks within both the branch and trunk networks of the DeepONet architecture. This integration allows the model to capture prediction uncertainties, enhancing its robustness against overfitting. However, VB-DeepONet faces challenges related to posterior inference using standard variational inference. A common practice in VB-DeepONet is to employ a fully factorised standard Gaussian distribution as the prior for the DeepONet model parameter vector $\bm{{\theta}}$. This prior assumes pairwise independence among model parameters, implying that each parameter distribution is unimodal. Although intended to be non-informative, this choice of prior can still exert significant influence on the posterior distribution given the number of parameters of deep neural networks can be quite large. Hence, this prior might not accurately reflect the true distribution of the parameters, potentially leading to sub-optimal uncertainty quantification. Moreover, standard KLD-VI with mean-field Gaussian variational families often exhibit “mode-seeking” behaviour, where the approximated posterior tends to concentrate around a single point. Additionally, if the model parameters exhibit high correlations, these VI approximations can produce overly confident predictions, failing to capture the true range of uncertainty. These issues can significantly hinder the VB-DeepONet’s ability to effectively quantify uncertainties, a crucial aspect for reliable engineering applications.
The next section discusses the proposed $\alpha$-VI DeepONet approach that uses an alternative divergence measure, the Rényi’s-$\alpha$ divergence, to address the limitations of VI in VB-DeepONet and achieve robust uncertainty quantification.

### v2

@@ Abstract
We introduce a novel deep operator network (DeepONet) framework that incorporates generalized variational inference (GVI) using Rényi’s $\alpha$-divergence to learn complex operators while quantifying uncertainty. By incorporating Bayesian neural networks as the building blocks for the branch and trunk networks, our framework endows DeepONet with uncertainty quantification. The use of Rényi’s $\alpha$-divergence, instead of the Kullback-Leibler divergence (KLD), commonly used in standard variational inference (VI), mitigates issues related to prior misspecification that are prevalent in Variational Bayesian DeepONets. This approach offers enhanced flexibility and robustness. We demonstrate that modifying the variational objective function yields superior results in terms of minimising the mean squared error and improving the negative log-likelihood on the test set. Our framework’s efficacy is validated across various mechanical systems, where it outperforms both deterministic and standard KLD-based VI DeepONets in predictive accuracy and uncertainty quantification. The hyperparameter $\alpha$, which controls the degree of robustness, can be tuned to optimise performance for specific problems. We apply this approach to a range of mechanics problems, including gravity pendulum, advection-diffusion, and diffusion-reaction systems. Our findings underscore the potential of $\alpha$-VI DeepONet to advance the field of data-driven operator learning and its applications in engineering and scientific domains.
Keywords Prior-robust Bayesian inference $\cdot$ Variational Bayes DeepONet $\cdot$ Bayesian neural networks $\cdot$ Bayesian inference $\cdot$ Uncertainty quantification $\cdot$ PDE surrogate
@@ 1 Introduction
As scientific machine learning methodologies take centre stage across diverse disciplines in science and engineering, there is an increased interest in adopting data-driven methods to analyse, emulate, and optimise complex physical systems. The behaviour of such systems is often described by laws expressed as systems of ordinary differential equations (ODEs) and partial differential equations (PDEs) [courant2008methods]. A classical task then involves the use of analytical or computational tools to solve such equations across a range of scenarios, e.g., different domain geometries, input parameters, and initial and boundary conditions (IBCs). Solving these so-called parametric PDE problems requires learning the solution operator that maps variable input entities to the corresponding solution of the underlying PDE system. Tackling this task using traditional tools (e.g., finite element methods [hughes2012finite]) bears a formidable cost, as independent simulations need to be performed for every different domain geometry, input parameters, or IBCs. Driven by this challenge, a growing field of neural operators for solving parametric PDEs has come forth recently.
@@
Learning to solve PDEs is closely related to operator learning. Instead of learning to solve a specific PDE, it can be advantageous to learn the operator that maps a functional parameter of the PDE (such as initial values, boundary conditions, force fields, or material parameters) to the solution associated with the given parameter. Neural operators exemplify this approach by learning to solve entire classes of PDEs simultaneously. These specially designed deep neural networks can represent the solution map of parametric PDEs in a discretisation-invariant manner, meaning the model can be queried at any arbitrary output location. Unlike neural networks that map between two finite-dimensional vector spaces, neural operators can represent mappings between spaces with infinite dimensions [kovachki2023neural]. Popular operator learning frameworks include deep operator networks (DeepONets) [lu2021learning] and Fourier neural operator (FNO) [li2020fourier]. This article focuses on DeepONet, which has been applied to a wide range of problems, including solid and fluid mechanics [goswami2023physics], reliability analysis [garg2022assessment], heat transfer [lu2022multifidelity], and fracture mechanics [goswami2022physics]. De et al. [de2022bi] utilised DeepONets for modelling uncertain and partially unknown systems. Cai et al. [cai2021deepm] applied DeepONets to model field variables across multiple scales in multi-physics problems, while He et al. [he2023novel] employed DeepONets to predict full-field, non-linear elastic–plastic stress responses in complex geometries. Other notable applications of DeepONets include weather forecasting [pathak2022fourcastnet], reduced-order modelling [demo2023deeponet], finance [leite2021deeponets], and others [ranade2021generalized, xu2024multi].
@@
While DeepONets excel at solving differential equations (DEs), a crucial aspect often overlooked is the uncertainty quantification of the predictions. Traditional DeepONet architectures have been found to produce over-confident predictions, implying they are poorly calibrated [lu2022comprehensive]. This means they might underestimate the true range of possible outcomes, potentially leading to unreliable results. Additionally, they do not quantify the inherent uncertainties associated with their predictions. In engineering applications like aerodynamics or structural analysis, uncertainty quantification is paramount. By quantifying uncertainties in predictions related to phenomena such as stress distribution, engineers can make informed decisions about the risk and reliability of their designs.
@@
Several attempts have been made to incorporate uncertainty quantification within the DeepONet architecture. However, existing approaches face some limitations. Lin et al. [lin2021accelerated] proposed a Bayesian DeepONet based on replica-exchange stochastic gradient Langevin diffusion [welling2011bayesian] that considers the standard deviation of the output to be known, which may not always be realistic. Additionally, training the replicas in this approach is computationally expensive. Yang et al. [yang2022scalable] used randomised priors and trained an ensemble of deterministic models, which can also be computationally demanding depending upon the size of the ensemble. While the method of ensembles offers some degree of uncertainty estimation through multiple model predictions, it does not adhere to the more principled framework that is provided by Bayes’ rule and hence cannot be considered truly Bayesian.
@@
Variational inference (VI)-based Bayesian methods offer a promising framework for incorporating uncertainty quantification into deep learning models. This approach has given rise to Bayesian neural networks (BNNs) [mackay1992practical, neal2012bayesian]. In BNNs, the parameters of the neural network are treated as random variables, which allows the uncertainty in the parameters to propagate through the network, ultimately reflecting the uncertainty in the model predictions. The variational inference approach is particularly advantageous for neural networks with a large number of parameters, as traditional Markov chain Monte Carlo (MCMC)-based methods become computationally expensive to use. Variational inference, on the other hand, provides a scalable alternative for approximating the posterior distribution of the parameters in BNN.
@@
It seems natural to extend the idea of uncertainty quantification in BNNs to that of a Bayesian DeepONet. Building on this idea, Garg et al. [garg2023vb] proposed VB-DeepONet that uses BNNs as building blocks for the DeepONet architecture. However, VB-DeepONet inherits certain challenges associated with BNNs. Their approach employs a fully factorised standard Gaussian distribution as the prior over the DeepONet model parameters. Such a prior might not accurately reflect the true distribution of these parameters, leading to a sub-optimal approximation of the posterior distribution. As indicated in Li et al. [li2016renyi], Knoblauch et al. [knoblauch2019generalized], and Wenzel et al. [wenzel2020good], such a misspecified prior can hinder the overall effectiveness of variational inference, where the approximated posterior concentrates around a single point, leading to overconfident predictions and underestimated uncertainty [Turner_Sahani_2011]. This hinders the model’s ability to effectively quantify uncertainties. Recent studies in Bayesian deep learning highlight that commonly used priors, such as isotropic Gaussians, can be unintentionally informative and lead to the so-called cold posterior effect [wenzel2020good]. Fortuin et al. [fortuin2021bayesian] further showed that isotropic Gaussian priors are often suboptimal for Bayesian neural networks and that exploring alternative priors is beneficial, especially as model depth and capacity increase. These works suggest that prior misspecification is not only a theoretical concern but has been empirically observed in practice, motivating the development of both more expressive priors and robust inference procedures.
@@
We propose to address these limitations in VB-DeepONet by introducing prior-robust variational inference for DeepONets. Standard VI relies on the Kullback-Leibler divergence (KLD), which is sensitive to prior selection, especially when the number of parameters is large. Alternative divergence measures exist which offer robustness to misspecified priors. Knoblauch et al. [knoblauch2019generalized] introduced the Generalized Variational Inference (GVI) framework that allows for defining divergence metrics beyond KLD in the context of BNNs. GVI incorporates a hyperparameter that controls the degree of robustness to prior misspecification. This leads to posteriors that are less influenced by priors which deviate significantly from the observed data [knoblauch2019generalized].
@@
In this paper, we extend the GVI framework to DeepONets to achieve greater robustness to prior misspecification. In particular, we propose the use of Rényi’s $\alpha$-divergence [renyi1961measures] as a robust alternative to KLD within the VI framework. This novel operator learning approach using GVI leads to an improved approximated posterior distribution, demonstrably enhancing both prediction accuracy and the quality of uncertainty estimates as evidenced by log-likelihood values on unseen test data. The rest of the article will provide background on deterministic DeepONets, followed by details of the proposed variational Bayesian framework and four numerical examples highlighting the efficacy of our approach. We conclude by discussing the results of different hyperparameter settings and recovering standard KLD-VI results as a special case.
@@ 2 Background on DeepONets
Operator networks were first introduced by Chen and Chen [chen1995universal], where they considered shallow networks with a single hidden layer. Subsequently, this concept was significantly developed with deep neural network architectures, leading to the creation of DeepONets [lu2021learning]. To define operators, it is useful to consider two different classes of functions residing in two separate Banach spaces, $\mathcal{A}$ and $\mathcal{S}$, respectively. One class of functions, $a(\bm{{y}})\in\mathcal{A}$, with domain in $\Omega_{Y}\in\mathbb{R}^{D}$, while the other class of functions, $s(\bm{{y}})\in\mathcal{S}$, also having domain in $\Omega_{Y}$. An operator $\mathcal{G}$ can be defined as a mapping between these two function spaces, $\mathcal{G}:\mathcal{A}\rightarrow\mathcal{S}$ such that $s(\bm{{y}})=\mathcal{G}(a(\bm{{y}}))$.
@@ 2.1 Deterministic DeepONets
Here, $\mathcal{F}$ is a non-linear differential operator, $x$ is the location in space, $t$ is the time; $s$ denotes the solution of the differential equation, and $u$ denotes the external source function. Furthermore, $\mathcal{B}(s)(x,t)=s_{b}(x)$ and $\mathcal{I}(s)(x,0)=s_{0}(x)$ define the boundary and initial conditions, respectively. The true solution operator is denoted by $\mathcal{G}(u)(x,t)$. To keep the notations compact, we denote the tuple $(x,t)$ by $\bm{{y}}$.
DeepONets [lu2021learning] are designed to learn an approximate operator ${\mathcal{G}}_{\theta}$ parameterised by a deep neural network with parameter vector $\bm{{\theta}}$.
@@
The architecture for the DeepONet consists of two main sub-networks: a branch neural network and a trunk neural network (see Fig.˜1).
@@
$\begin{split}\hat{\bm{{\theta}}}&=\operatorname*{argmin}_{\bm{{\theta}}}\frac{1}{N_{1}N_{2}}\sum_{i=1}^{N_{1}}\sum_{k=1}^{N_{2}}\left(\mathcal{G}_{\bm{{\theta}}}\left(\bm{{a}}^{(i)}\right)(\bm{{y}}_{k})-s^{(i)}\left(\bm{{y}}_{k}\right)\right)^{2}\\ &=\operatorname*{argmin}_{\bm{{\theta}}_{B},\bm{{\theta}}_{T}}\frac{1}{N_{1}N_{2}}\sum_{i=1}^{N_{1}}\sum_{k=1}^{N_{2}}\left(b_{0}+\sum_{p=1}^{P}b_{p}\left(\bm{{a}}^{(i)};\bm{{\theta}}_{B}\right)\;\psi_{p}\left(\bm{{y}}_{k};\bm{{\theta}}_{T}\right)-s^{(i)}\left(\bm{{y}}_{k}\right)\right)^{2}.\end{split}$ (4)
The optimisation problem can be solved using stochastic gradient descent [ruder2016overview] to obtain the optimal parameter values for the DeepONet. This approach is referred to as D-DeepONet in the rest of the article, where “D” stands for deterministic.
@@
2.2 Uncertainty quantification in DeepONets
Several attempts have been made to incorporate uncertainty quantification within the DeepONet architecture. Early efforts focused on Bayesian neural network formulations, such as replica-exchange stochastic gradient Langevin diffusion [lin2021accelerated] or randomised-prior ensembles [yang2022scalable], as well as baselines like classical ensembles and Monte Carlo Dropout [psaros2023uncertainty].
Beyond these BNN-based methods, several alternative frameworks have emerged. Conformalized-DeepONet applies split conformal prediction (including a Quantile-DeepONet variant), yielding distribution-free prediction intervals with finite-sample coverage guarantees, albeit requiring a calibration split and sometimes producing conservative intervals under strong distribution shift [moya2025conformalized]. The Information Bottleneck-UQ framework introduces an information-bottleneck objective with a confidence-aware encoder and Gaussian decoder, offering computationally efficient and out-of-distribution aware UQ for DeepONet [guo2024ib]. Outside neural parameterisations, kernel and Gaussian-process based operator learning provide mathematically interpretable priors and exact GP-style uncertainty estimates; recent results show kernel methods are competitive with, and sometimes outperform, neural operators on benchmark tasks, while GP-NN hybrids combine exact GP-based UQ with flexible neural mean functions [mora2025operator, batlle2024kernel]. Another promising direction is Ensemble Kalman Inversion for DeepONet [pensoneault2025uncertainty], which delivers practical epistemic UQ with strong parallelisation, though at the expense of additional computational cost and hyperparameter tuning.
Taken together, these methods extend the UQ toolbox for operator learning beyond Bayesian neural networks, striking different balances between rigour, scalability, and interpretability.
@@ 2.3 Variational Bayesian DeepONets (VB-DeepONet)
VB-DeepONet [garg2023vb] addresses the limitations of D-DeepONet by incorporating layers of Bayesian neural networks within both the branch and trunk networks of the DeepONet architecture. This integration allows the model to capture prediction uncertainties, enhancing its robustness against overfitting. However, VB-DeepONet faces challenges related to posterior inference using standard variational inference. A common practice in VB-DeepONet is to employ a fully factorised standard Gaussian distribution as the prior for the DeepONet model parameter vector $\bm{{\theta}}$. This prior assumes pairwise independence among model parameters, implying that each parameter distribution is unimodal. Although intended to be non-informative, this choice of prior can still exert significant influence on the posterior distribution, given that the number of parameters of deep neural networks can be quite large. Hence, this prior might not accurately reflect the true distribution of the parameters, potentially leading to sub-optimal uncertainty quantification. Moreover, standard KLD-VI with mean-field Gaussian variational families often exhibit “mode-seeking” behaviour, where the approximated posterior tends to concentrate around a single point. Additionally, if the model parameters exhibit high correlations, these VI approximations can produce overly confident predictions, failing to capture the true range of uncertainty. These issues can significantly hinder the VB-DeepONet’s ability to effectively quantify uncertainties, a crucial aspect for reliable engineering applications.
The next section discusses the proposed $\alpha$-VI DeepONet approach that uses an alternative divergence measure, the Rényi’s $\alpha$-divergence, to address the limitations of VI in VB-DeepONet and achieve robust uncertainty quantification.